"Find an angle theta in quadrant IV where sin(theta)=sin(8pie/7)"
Answer by stanbon(75887)

Find an angle theta in quadrant IV where sin(theta)=sin(8pie/7)"
-----
Reference angle = pi/7
-----
(8/7)pi is in QIII
----
In QIV sin(13pi/7) = sin(8pi/7)
----
Ans: theta = 13pi/7
----------------
Cheers,
Stan H.